Determine if the following functions are continuous or discontinous at the given value of x.



1. f(x)= 3x+2x+1 at x= -2


2. f(x)= 9x2-1 at x=1


3. f(x)= 1/x-2 at x=2


4. f(x)= x-1/x^2 -1 at x=1


5. h(x)= x+1/x-1at x=1


6. f(x)= x^2 -2x+1 at x=3


7. f(x)= x/(x+2)(x-3) at x=2


8. f(x)= |x| at x=0


9. f(x) Cube root√x^2 -4 at x=5


10. f(x)= x+1/x-1 at x=0



1
Expert's answer
2022-02-22T14:58:26-0500

1. Since "any polynomial function is continuous everywhere", "f(x) = 3x^2+2x+1" is continuous everywhere and hence continuous at x = -2


2. Since "f(x)= 9x^2-1" is a polynomial, it is continuous at x = 1.


3. The function "f(x)= \dfrac{1}{x-2}" is discontinuous at x = 2 and it is an infinite discontinuity.


4. For the function "f(x)= \dfrac{x-1}{x^2 -1}", since both the numerator and the denominator are continuous, the discontinuities are the points obtained from "x^2-1=0\text{~i.e.,~} x = \pm1", where "x = 1" is a removable discontinuity since, "f(x)= \dfrac{x-1}{x^2 -1} = \dfrac{1}{x+1}"becomes continuous at "x = 1".


5. The function "h(x)= \dfrac{x+1}{x-1}" is discontinuous at x = 1, since "\displaystyle\lim_{x \rightarrow 1} h(x)\rightarrow\infty".


6. The function "f(x)= x^2 -2x+1" is continuous at x=3 as it is a polynomial function.


7. Since both numerator and denominator of the function "f(x)= \dfrac{x}{(x+2)(x-3)}" are continuous, the points of discontinuity are, "(x+2)(x-3)=0", i.e., "x = -2,3". Therefore, "f(x)" is continuous at "x = 2".


8. "f(x)= |x| =\begin{cases}-x, &x <0\\ x, &x\geq 0\end{cases}".

We know that f(x) is continuous at x = 0 if "\displaystyle\lim _{x \rightarrow 0^{-}} f(x) = \lim _{x \rightarrow 0^{+}} f(x) = f(0)".

Now,

"\begin{aligned}
\lim _{x \rightarrow 0^{-}} f(x) &=\lim _{h \rightarrow 0} f(0-h) \\
&=\lim _{h \rightarrow 0} f(-h) \\
&=\lim _{h \rightarrow 0}|-h| \\
&=\lim _{h \rightarrow 0} h \\
&=0
\end{aligned}"

"\begin{aligned}
\lim _{x \rightarrow 0^{+}} f(x) &=\lim _{h \rightarrow 0} f(0+h) \\
&=\lim _{h \rightarrow 0} f(h) \\
&=\lim _{h \rightarrow 0}|h| \\
&=\lim _{h \rightarrow 0} h \\
&=0
\end{aligned}"

and "f(0) = 0". Hence f(x) is continuous at x = 0.


9. Since "x^2 - 4" is a polynomial which is continuous everywhere, "f(x)= \sqrt[3]{x^2 -4}" is continuous at x = 5.


10. For the function "f(x)= \dfrac{x+1}{x-1}" the point of discontinuity is x= 1, and hence f(x) is continuous at x = 0.